### v2

@@ 3.3 Retrain scenarios
To answer our research question, we explored several possible retraining scenarios. A retrain scenario or retrain window, $r$, is identified as a positive integer representing the frequency at which the model is re-trained or updated. Essentially, it represents how many data points need to be passed before a new training step is performed. The retrain scenarios strongly depend on the frequency of the dataset considered because the frequency drives both the forecast horizon and the business review periods. Therefore, we defined different set of retrain scenarios based on the frequencies of our datasets:
  • 1.
    for weekly data, $r=\{1,2,3,4,6,8,10,13,26,52\}$
  • 2.
    for daily data, $r=\{7,14,21,30,60,90,120,150,180,364\}$
@@ 3.4 Evaluation setup
In time series forecasting, out-of-sample testing is crucial for assessing how well a model can generalize to unseen data. This is particularly important because patterns in historical data may not fully represent future trends, and unexpected changes in the data may occur (Tashman, 2000). As suggested by Bergmeir & Benítez (2012), rolling origin evaluation is the most widely used and correct method for conducting such tests, offering a systematic way to evaluate forecasting models over multiple iterations. The rolling origin evaluation process begins by dividing the time series into two parts: a training set, which is used to build the model, and a test set, which is used for evaluation. The test set always follows the training set chronologically to maintain the temporal order of the data. At each step, the model is trained on the training set and generates forecasts for a specified number of future points, the forecast horizon $h$. These forecasts are then compared to the actual values in the test set using a chosen evaluation metric (see Section 3.5). What makes rolling origin evaluation unique is its iterative nature. After each forecasting step, the forecast origin (the last point in the training set) is shifted forward by a fixed number of periods, the step size. The model is then retrained using the updated training set, which may either expand to include all available data up to the new forecast origin or remain fixed to a specific window size. This process is repeated until the test set is fully utilized, and the overall performance of the model is calculated by averaging the results across all the iterations.
@@
Compared to the fixed origin evaluation setup, which allows for a single evaluation step, the rolling origin evaluation is preferable for its ability to provide a more comprehensive assessment of a model’s performance under different conditions, such as varying seasonal patterns, level shifts, or data trends. By simulating multiple forecast cycles, indeed, this method reduces the risk of bias that can arise from evaluating a model a single time and at a single forecast origin (Bergmeir & Benítez, 2012). This approach is beneficial in industries like retail and supply chain management, where forecasts need to be continuously updated to reflect real-time changes. By simulating multiple scenarios and evaluating a model’s performance across different cycles, rolling origin evaluation provides valuable insights into the robustness and reliability of the forecasting methods adopted.
@@
In practice, the rolling origin evaluation setup can be adjusted based on the data and forecasting goals. For instance, the training set can be fixed to a rolling window of the most $n$ recent observations (fixed window), or it can expand to include the full historical data (expanding window) (Bergmeir & Benítez, 2012). In real-life applications, most forecast practitioners use this second option, especially when the length of the time series is small (Petropoulos & et al., 2022). In our study, we adopted the expanding window approach because we are interested in simulating a forecasting experiment as close as possible to the business reality. Moreover, it is also the only reasonable option when dealing with short time series, like the case for the weekly VN1 datasets.
Table 2: Retraining scenarios, test windows, and forecast horizons for each dataset. The setup is based on the respective time series frequency.
M5 VN1
Training Window ($w$) $\geq 730$ $\geq 157$
Test Window ($T$) 364 52
Horizon ($h$) 28 13
Step size ($p$) 1 1
Retraining Scenarios ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
@@
The size of the test set $T$, the length of the forecast horizon $h$, and the step size $p$, for shifting the forecast origin, are also customizable parameters that depend on the specific use case. Table 2 shows the parameters selected in our experiments. Following Spiliotis & Petropoulos (2024), we set the size of the test set to cover at least one complete year so that the evaluation of the different scenarios does not depend on intra-year variations, such as any specific season or period of the year. We selected the forecast horizons based on the type of business decisions that usually daily and weekly forecasts support, i.e. operational vs strategic planning. Finally, the step size was always set equal to one to maximize the number of evaluations of each scenario.
@@ 3.5 Performance metrics
The accuracy evaluation of point predictive models is a controversial topic in time series forecasting: many metrics are available to capture models’ performances, but no consensus has been reached in the literature on whether one metric is better than others (Hewamalage et al., 2023). However, since we are dealing with demand forecasting at SKU-level, implying that data intermittency is very much likely, all the metrics based on absolute or percentage errors are not optimal since they optimize for the median (Kolassa, 2020). Moreover, being the scale possibly different among the series, a scaled accuracy measure is preferable. For this reason, we considered the Root Mean Squared Scaled Error (RMSSE), proposed by Hyndman & Koehler (2006), to evaluate the point forecast accuracy of the models. The RMSSE is defined as:
@@
The RMSSE measures the relative prediction accuracy of a forecasting method by comparing the mean squared errors of the prediction and the observed value against the mean squared errors of the seasonal naive model. This metric was the official metric used to assess the performance of models in the M5 competition (with $s=1$, hence using the naive model as a benchmark) (Makridakis et al., 2022b). Lower RMSSE values suggest higher model accuracy.
Our work is not limited to the evaluation of point forecasting accuracy, but we also aim to assess the probabilistic performance of the models throughout each different retrain scenario. However, since the machine learning and deep learning models adopted for the experiments do not output density (or probabilistic) predictions by default, we relied on Conformal Inference as a general framework to build prediction intervals. Conformal Inference is a powerful tool that allows quantification of the uncertainty associated with out-of-sample predictions simply using a validation set on a point forecaster model to produce intervals at the desired levels (Vovk et al., 2005). It was originally proposed for non-temporal data, because the data exchangeability assumption, nevertheless, in recent years it has been extended to be used also in the context of time series data (Stankeviciute et al., 2021). In particular, given its guaranteed coverage, distribution-free, model-agnostic, computational efficiency, and low dataset size requirements, Conformal Inference is likely the only tool available to evaluate the uncertainty in a global forecasting setting where different models are compared among several datasets and scenarios. One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, but this would require training a single model for each quantile. Even if it may be a valuable way of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022c; Lainder & Wolfinger, 2022), it has the huge disadvantage of being very computationally expensive, hence negatively impacting on the sustainability of forecasting applications in large-scale environments. Moreover, even if multi-quantile losses can provide multiple quantiles, and prediction intervals, they usually are uncalibrated, meaning that the actual frequency of observations falling within the interval does not align with the confidence level associated with it. Therefore, for these multiple reasons, we decided to adopt Conformal Inference as general tool to estimate quantile forecasts.
Once the quantiles are obtained, we used a scaled version of the Multi-Quantile Loss to comprehensively measure the goodness of the probabilistic predictions. The scaled Quantile Loss, also known as (scaled) Pinball Loss, and the scaled Multi-Quantile Loss are defined as:
@@
where $\mathcal{Q}$ is the set of possible quantiles. The QL, being a proper scoring rule, allows for the correct evaluation of the probabilistic forecasts (Kolassa, 2016). The scaling factor is the in-sample, one-step-ahead mean absolute error of the seasonal Naive model. The scaled MQL (weighted and with $s=1$) was also the official metric used to evaluate the overall performance of the forecasting methods during the M5 Uncertainty competition (Makridakis et al., 2022c). For daily data we used $s=7$, because $s=1$ is arguably an irrelevant benchmark in daily retail data, which typically exhibits strong day-of-week seasonality (Fildes et al., 2022). Instead, we set $s=1$ for weekly data.
Note that we considered a total of 14 different quantiles, $q$, of probability levels {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we studied the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, which in retail demand forecasting problems are fundamental to determine the appropriate safety stock levels (Barrow & Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
Note also that we computed conformal prediction intervals on a validation set at least twice as long as the forecast horizon. This was the main reason that forced us to reduce the number of time series in the datasets, that is to be able to obtain a reliable estimate of the model uncertainty through prediction intervals 777 We used a validation set four times larger than the forecast horizon for the M5 dataset and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 30,490 to 28,298 for the M5 dataset. .
Finally, we addressed the problem of evaluating the cost associated to producing the forecasts by calculating the Computing Time (CT). CT is defined as the time in seconds required for training and predicting the next $h$ time steps ahead with the model. Among the possible measures of complexity, such as the number of parameters, number of iterations, model’s depth, etc, CT is the most directly related to the monetary cost of forecasting since the forecasts are usually produced through some sort of cloud computing service that relies on pay-as-you-go fares (Spiliotis & Petropoulos, 2024). Hence, CT is analyzed for each forecasting model and retraining scenario combination. Lower CT values imply lower forecasting costs.
@@ 4 Results and discussion
Figure 2 statistically confirms the above results for the M5 dataset 888 The Friedman-Nemenyi test is usually adopted to compare and rank the performance of different models. Here we use it to compare the accuracy produced by the same model over different retraining scenarios. . It is clear that the retraining scenarios are not statistically different in terms of point forecast accuracy at the 5% level. Indeed, even though the mean rank of the continuous retraining scenario is lower, the intervals intersect with the other scenarios, hence we can not differentiate between them. For the VN1 dataset (see Supplementary materials), less frequent retrainings are even statistically significant, meaning that periodic retraining does improve accuracy over continuous retraining.
@@
Figure 2: M5 Friedman-Nemenyi test results based on RMSSE.
@@
The statistical test to compare probabilistic forecasting performances among the different retraining scenarios on the M5 dataset, as depicted in Figure 4, confirms that there are significant differences in accuracy. In particular, lower restraining scenarios tend to produce more accurate probabilistic forecasts, that is, less frequent retraining effectively harms the accuracy. However, this is not true for low retraining scenarios where the differences are not statistically significant, implying that even in the context of probabilistic forecasting the frequency of retraining may be reduced to at least once a month (or every two months). For the VN1 (results in Supplementary materials), instead, some level of periodic retaining is better than continuous retraining, most of the time.
@@
Figure 4: M5 Friedman-Nemenyi test results based on SMQL.
@@
Overall, the results of Figures 1, 3, and 5, combined with the Friedman-Nemenyi tests, suggest that retraining the global models less frequently does not harm (or even improves) forecast accuracy, while at the same time significantly reducing the computing time of producing the forecasts. Thus, extending the retraining time from the standard practice of continuous retaining to some level of periodic retraining allows for effective management of the computing time, and in turn, the costs of forecasting. Indeed, as already observed, the computational time can be directly translated into actual costs for the company. Following Nikolopoulos & Petropoulos (2018) and Fotios Petropoulos & Spiliotis (2024), we assumed some standard costs for computing services to estimate the costs of forecasting associated with each retrain scenario. Note that costs and savings are normalized by the number of series considered in each dataset (28,298 and 15,053) so that they can be directly compared and conclusions can be drawn in terms of the frequency of the time series. Figures 6, and 7 show the costs and associated savings for a large retailer, given a fixed computing service cost of \$3.5/hour, 200,000 unique SKUs and 5,000 stores, which is approximately the size of the forecasting problem of Walmart (Spiliotis & Petropoulos, 2024). As expected, the costs decrease exponentially with less frequent retraining. For daily data, forecasting with machine learning models usually costs less than with deep learning models. On average, the continuous retraining scenario costs approximately \$750,000 and this cost drops down to almost \$250,000 in the no retraining scenario, implying direct savings of more than 60%. Moreover, machine learning models allow for higher savings. Indeed, even if the models have to be updated, going from retraining every week to retraining every month allows to get direct savings of almost 75%, while it can be less than 30% for deep learning models. This implies that machine learning methods are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining as the frequency of the data increases. For weekly data, these differences are less evident, meaning that machine learning and deep learning models usually have very similar costs and savings profiles (without considering the Random Forest model, which is usually more than 10 times slower than other methods). This is to be expected since lower frequencies also imply smaller datasets and lower computing time. In this case, the average cost of forecasting under a continuous retraining scenario is \$250,000 (1/3 of that of daily data), and it drops to \$15,000 in the least frequent retraining scenario, producing direct savings of 90%. Nevertheless, for weekly data too, moving from retraining a model every week ($r=1$) to just once a month ($r=4$) allows to reduce the cost of almost 75%.
@@
Finally, in this study, since we are dealing with global models, we assumed a fixed model retraining frequency across all forecast origins and time series. While our results showed no significant difference in forecasting accuracy when reducing the retraining frequency, we believe there is potential to further improve accuracy, especially when adopting some form of periodic retraining, for instance, by allowing the retraining frequency to vary dynamically. Such adaptive strategies could involve automated procedures that monitor forecast error levels and trigger updates only when errors exceed a certain threshold. This approach would still yield computational savings, albeit smaller, since the model re-training would not be required at every review point. In this direction, Figures 8 and 9 illustrate the ”optimal” retraining frequency identified for each series in the M5 and VN1 datasets, considering all the models, based on our experiments. The results confirm that although certain series may benefit from frequent retraining, in most cases, a less frequent retraining of the models leads to better accuracy and efficiency. Indeed, for point forecasting, the optimal retraining frequency was found to be between 3 and 4 weeks (approximately 21 observations) for the M5 dataset, and around 8 to 10 weeks for the VN1 dataset. In the context of probabilistic forecasting, instead, the optimal retraining intervals were shorter (around 2 weeks for M5 and 4 weeks for VN1), suggesting that uncertainty estimation may require more frequent updates to maintain accurate levels. Moreover, while these results do not differentiate among the global forecasting methods used in the experiment, individual model differences exist, implying that even further improvements may be obtained by relying on specific model (see Supplementary material). These findings suggest that exploring intermediate retraining strategies and determining the optimal retraining frequency offers promising avenues for reducing the computational cost of the forecasting process while maintaining a high level of accuracy.
@@
Figure 9: Overall optimal retraining frequency based on SMQL. For each series, the optimal frequency is identified based on the average accuracy reported across the evaluation rounds per update frequency, considering all the models.

### v3

@@ 3.3 Retrain scenarios
To answer our research question, we explored several possible retraining scenarios. A retrain scenario or retrain window, $r$, is identified as a positive integer representing the frequency at which the model is re-trained or updated. Essentially, it represents how many data points need to be passed before a new training step is performed. The retrain scenarios strongly depend on the frequency of the dataset considered because the frequency drives both the forecast horizon and the business review periods. Therefore, we defined two different sets of retraining scenarios based on the frequency of the datasets.
Table 1: Retraining set, train and test size, frequency, forecast horizon, and step size for each dataset. The setup is based on the respective time series frequency.
M5 VN1
Frequency ($f$) daily (7) weekly (52)
Train size ($n$) $\geq 730$ $\geq 157$
Test size ($T$) 364 52
Horizon ($h$) 28 13
Step size ($p$) 1 1
Retrain set ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
@@ 3.4 Evaluation setup
In time series forecasting, out-of-sample testing is crucial for assessing how well a model can generalize to unseen data. This is particularly important because patterns in historical data may not fully represent future trends, and unexpected changes in the data may occur (Tashman, 2000). As suggested by Bergmeir and Benítez (2012), rolling origin evaluation is the most widely used and correct method for conducting such tests, offering a systematic way to evaluate forecasting models over multiple iterations. The rolling origin evaluation process begins by dividing the time series into two parts: a training set, which is used to build the model, and a test set, which is used for evaluation. The test set always follows the training set chronologically to maintain the temporal order of the data. At each step, the model is trained on the training set and generates forecasts for a specified number of future points, the forecast horizon $h$. These forecasts are then compared to the actual values in the test set using a chosen evaluation metric (see Section 3.5). What makes rolling origin evaluation unique is its iterative nature. After each forecasting step, the forecast origin (the last point in the training set) is shifted forward by a fixed number of periods, the step size. The model is then retrained using the updated training set, which may either expand to include all available data up to the new forecast origin or remain fixed to a specific window size. This process is repeated until the test set is fully utilized, and the overall performance of the model is calculated by averaging the results across all the iterations.
@@
Compared to the fixed origin evaluation setup, which allows for a single evaluation step, the rolling origin evaluation is preferable for its ability to provide a more comprehensive assessment of a model’s performance under different conditions, such as varying seasonal patterns, level shifts, or data trends. By simulating multiple forecast cycles, indeed, this method reduces the risk of bias that can arise from evaluating a model a single time and at a single forecast origin (Bergmeir and Benítez, 2012). This approach is beneficial in industries like retail and supply chain management, where forecasts need to be continuously updated to reflect real-time changes. By simulating multiple scenarios and evaluating a model’s performance across different cycles, rolling origin evaluation provides valuable insights into the robustness and reliability of the forecasting methods adopted.
@@
In practice, the rolling origin evaluation setup can be adjusted based on the data and forecasting goals. For instance, the training set can be fixed to a rolling window of the most $n$ recent observations (fixed window), or it can expand to include the full historical data (expanding window) (Bergmeir and Benítez, 2012). In real-life applications, most forecast practitioners use this second option, especially when the length of the time series is small (Petropoulos and et al., 2022). In our study, we adopted the expanding window approach because we are interested in simulating a forecasting experiment as close as possible to the business reality. Moreover, it is also the only reasonable option when dealing with short time series, like the case for the weekly VN1 datasets.
@@
The size of the test set $T$, the length of the forecast horizon $h$, and the step size $p$, for shifting the forecast origin, are also customizable parameters that depend on the specific use case. Table 1 shows the parameters selected in our experiments. Following Spiliotis and Petropoulos (2024), we set the size of the test set to cover at least one complete year so that the evaluation of the different scenarios does not depend on intra-year variations, such as any specific season or period of the year. We selected the forecast horizons based on the type of business decisions that usually daily and weekly forecasts support, i.e. operational vs strategic planning. Finally, the step size was always set equal to one to maximize the number of evaluations of each scenario.
@@ 3.5 Performance metrics
The accuracy evaluation of point predictive models is a controversial topic in time series forecasting: many metrics are available to capture models’ performances, but no consensus has been reached in the literature on whether one metric is better than others (Hewamalage et al., 2023). However, since we are dealing with demand forecasting at SKU-level, implying that data intermittency is very much likely, all the metrics based on absolute or percentage errors are not optimal since they optimize for the median (Kolassa, 2020). Moreover, being the scale possibly different among the series, a scaled accuracy measure is preferable. For this reason, we considered the Root Mean Squared Scaled Error (RMSSE), proposed by Hyndman and Koehler (2006), to evaluate the point forecast accuracy of the models. The RMSSE is defined as:
@@
The RMSSE measures the relative prediction accuracy of a forecasting method by comparing the mean squared errors of the prediction and the observed value against the mean squared errors of the seasonal naive model. This metric was the official metric used to assess the performance of models in the M5 competition (with $s=1$, hence using the naive model as a benchmark) (Makridakis et al., 2022c). Lower RMSSE values suggest higher model accuracy.
Our work is not limited to the evaluation of point forecasting accuracy, but we also aim to assess the probabilistic performance of the models throughout each different retrain scenario. We used a scaled version of the Multi-Quantile Loss to comprehensively measure the goodness of the probabilistic predictions. The scaled Quantile Loss, also known as (scaled) Pinball Loss, and the scaled Multi-Quantile Loss are defined as:
@@
where $\mathcal{Q}$ is the set of possible quantiles. The QL, being a proper scoring rule, allows for the correct evaluation of the probabilistic forecasts (Kolassa, 2016). The scaling factor is the in-sample, one-step-ahead mean absolute error of the seasonal Naive model. The scaled MQL (weighted and with $s=1$) was also the official metric used to evaluate the overall performance of the forecasting methods during the M5 Uncertainty competition (Makridakis et al., 2022a). For daily data we used $s=7$, because $s=1$ is arguably an irrelevant benchmark in daily retail data, which typically exhibits strong day-of-week seasonality (Fildes et al., 2022). Instead, we set $s=1$ for weekly data.
Note that we considered a total of 14 different quantiles, $q$, of probability levels {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.250, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we studied the 50% 60%, 70%, 80%, 90%, 95%, and 99% symmetric prediction intervals. The 50% and 60% symmetric prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails, which in retail demand forecasting problems are fundamental to determine the appropriate safety stock levels (Barrow and Kourentzes, 2016). These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution.
Finally, we addressed the problem of evaluating the cost associated to producing the forecasts by calculating the Computing Time (CT). CT is defined as the time in seconds required for training and predicting the next $h$ time steps ahead with the model. Among the possible measures of complexity, such as the number of parameters, number of iterations, model’s depth, etc, CT is the most directly related to the monetary cost of forecasting since the forecasts are usually produced through some sort of cloud computing service that relies on pay-as-you-go fares (Spiliotis and Petropoulos, 2024). Hence, CT is analyzed for each forecasting model and retraining scenario combination. Lower CT values imply lower forecasting costs.
@@ 4 Results and discussion
Overall, the results of Figures 1, 2, and 3, combined with the Friedman-Nemenyi tests for statistical significance (see Supplementary material), suggest that retraining the global models less frequently does not harm (or even improves) forecast accuracy, while at the same time significantly reducing the computing time of producing the forecasts. Thus, extending the retraining time from the standard practice of continuous retaining to some level of periodic retraining allows for effective management of the computing time, and in turn, the costs of forecasting. Indeed, as already observed, the computational time can be directly translated into actual costs for the company. Following Nikolopoulos and Petropoulos (2018) and Petropoulos and Spiliotis (2024), we assumed some standard costs for computing services to estimate the costs of forecasting associated with each retrain scenario. Note that costs and savings are normalized by the number of series considered in each dataset so that they can be directly compared and conclusions can be drawn in terms of the frequency of the time series. Figures 4, and 5 show the costs and associated savings for a large retailer, given a fixed computing service cost of \$3.5/hour, 200,000 unique SKUs and 5,000 stores, which is approximately the size of the forecasting problem of Walmart (Spiliotis and Petropoulos, 2024). As expected, the costs decrease exponentially with less frequent retraining. For daily data, forecasting with machine learning models usually costs less than with deep learning models. On average, the continuous retraining scenario costs approximately \$750,000 and this cost drops down to almost \$250,000 in the no retraining scenario, implying direct savings of more than 60%. Moreover, machine learning models allow for higher savings. Indeed, even if the models have to be updated, going from retraining every week to retraining every month allows to get direct savings of almost 75%, while it can be less than 30% for deep learning models. This implies that machine learning methods are a marginally better choice to reduce the costs of forecasting when coupled with less frequent model retraining as the frequency of the data increases. For weekly data, these differences are less evident, meaning that machine learning and deep learning models usually have very similar costs and savings profiles (without considering the Random Forest model, which is usually more than 10 times slower than other methods). This is to be expected since lower frequencies also imply smaller datasets and lower computing time. In this case, the average cost of forecasting under a continuous retraining scenario is \$250,000 (1/3 of that of daily data), and it drops to \$15,000 in the least frequent retraining scenario, producing direct savings of 90%. Nevertheless, for weekly data too, moving from retraining a model every week ($r=1$) to just once a month ($r=4$) allows to reduce the cost of almost 75%.
@@
Finally, in this study, since we are dealing with global models, we assumed a fixed model retraining frequency across all forecast origins and time series. While our results showed no significant difference in forecasting accuracy when reducing the retraining frequency, we believe there is potential to further improve accuracy, especially when adopting some form of periodic retraining, for instance, by allowing the retraining frequency to vary dynamically. Such adaptive strategies could involve automated procedures that monitor forecast error levels and trigger updates only when errors exceed a certain threshold. This approach would still yield computational savings, albeit smaller, since the model re-training would not be required at every review point. In this direction, Figure 6 illustrates the ”optimal” retraining frequency identified for each series in the M5 and VN1 datasets, considering all the models. The results confirm that although certain series may benefit from frequent retraining, in most cases, a less frequent retraining of the models leads to better accuracy and efficiency. Indeed, for point forecasting, the optimal retraining frequency was found to be between 3 and 4 weeks (approximately 21 observations) for the M5 dataset, and around 8 to 10 weeks for the VN1 dataset. In the context of probabilistic forecasting, instead, the optimal retraining intervals were shorter, around 2 weeks for M5 and 4 weeks for VN1 (see Supplementary material), suggesting that uncertainty estimation may require more frequent updates to maintain accurate levels. Moreover, while these results do not differentiate among the global forecasting methods used in the experiment, individual model differences exist, implying that even further improvements may be obtained by relying on specific model (see Supplementary material). These findings suggest that exploring intermediate retraining strategies and determining the optimal retraining frequency offers promising avenues for reducing the computational cost of the forecasting process while maintaining a high level of accuracy.